# Error Slice Discovery via Manifold Compactness

### v1

@@ A.5 Ablation of the slicing function
A.6 Choice of feature extractors
We conduct additional experiments on CelebA by changing CLIP-ViT-B/32 to CLIP-ResNet50 and ImageNet-supervised-pretrained ResNet50. Table 7 shows that whatever the pretrained feature extractor is, MCSD consistently identifies slices of low accuracy and outperforms other methods in terms of manifold compactness. It is worth noting that this conclusion is valid even for MCSD with ResNet50, which is generally considered as a weaker pretrained feature extractor than ViT-B/32 employed by baselines. In Figure 13, we can see that MCSD with different pretrained feature extractors truly identifies coherent error slices for the blond hair category of CelebA. As for the practice of using pretrained feature extractors, it is acceptable and generally adopted in previous works of error slice discovery (Eyuboglu et al., 2022; Jain et al., 2023; Plumb et al., 2023).
@@ A.6 Choice of feature extractors
Method Acc. (%) $\downarrow$ Comp.$\uparrow$ Acc. (%) $\downarrow$ Comp.$\uparrow$
@@
A.7 More examples for case studies
Figure 14: More examples of the category “Blond Hair” of CelebA.
Figure 15: More examples of the category “Not Blond Hair” of CelebA.
Figure 16: More examples of the category “ill” of CheXpert.
Figure 17: More examples of the category “healthy” of CheXpert.
A.8 Time comparison
@@
A.9 Hyperparameter selection and analyses
A.9.1 Hyperparameter selection
@@
A.9.2 Hyperparameter analyses
In this part, we conduct hyperparameter analyses on the category “Blond Hair” of CelebA for the coherence coefficient $\lambda$, the size $\alpha$, and the number of neighbors $k$ when building the kNN graph. From Table 9, we find that both the accuracy and manifold compactness are best when $\lambda$ and $\alpha$ are in a moderate range, neither too large nor too small. This implies the importance of the balance between pursuing high error and high coherence, which could be achieved by the tuning strategy mentioned in Section A.9.1. This also implies the importance of appropriately controlling $\alpha$, i.e. the size of the slice, which is set according to experience in our implementation. Its selection is left for future work.
@@ A.9.2 Hyperparameter analyses
$\lambda$ Acc. (%) $\downarrow$ Comp.$\uparrow$ $\alpha$ Acc. (%) $\downarrow$ Comp.$\uparrow$ $k$ Acc. (%) $\downarrow$ Comp.$\uparrow$
@@
A.10 Performance improvement via utilization of the discovered error slices
We conduct experiments to show performance improvement that the identified error slices could bring via data collection guided by the interpretable characteristics of identified error slices, following the practice of non-algorithmic interventions of Liu et al. (2023). For example, for a given trained image classification model on CelebA, the identified error slice exhibits characteristics of blond hair male, then we could be guided to collect specific data of the targeted characteristics of blond hair male and add to the training data, which is a non-algorithmic intervention and a straightforward and practical way of improving performance of the original model after interpreting characteristics of the identified error slice. Here we compare the results of guided data collection and random data collection. To simulate the guided data collection process, for CelebA, since the identified error slice for the category of blond hair is male and there are extra annotations of sex, we add the images annotated as blond hair male in validation data to training data. Since the identified error slice for not blond hair category is female bearing vintage styles, and there are no related attribute annotations, we directly add the images of the identified slice to the training data. For CheXpert, the identified error slices are from the frontal view for ill patients and from the left lateral view for healthy patients, and CheXpert has annotations of views, so we add the corresponding images in validation data to training data. To simulate the random data collection process, we randomly sample the same number of images from validation data and add to training data for each dataset. Then we retrain the model three times with varying random seeds.
@@ A.10 Performance improvement via utilization of the discovered error slices
Here worst group accuracy is defined following a distribution shift benchmark (Yang et al., 2023), where CelebA and CheXpert are divided into groups according to annotated attributes, and worst group accuracy is an important metric. From Table 10, we can see that guided data collection outperforms the original model and random data collection in both metrics, especially in worst group accuracy. This illustrates that our method is beneficial to performance improvement in practical applications.
A.11 Benchmark details
Dcbench (Eyuboglu et al., 2022) offers a large number of settings for the task of error slice discovery. Each setting consists of a trained ResNet-18 (He et al., 2016), a validation dataset and a test dataset, both with labels of predefined underperforming slices. The validation dataset and its error slice labels are taken as the input of slice discovery methods, while the test dataset and its error slice labels are used for evaluation.
There are 886 settings publicly available in the official repository of dcbench [^1] https://github.com/data-centric-ai/dcbench, comprising three types of slices: correlation slices, rare slices, and noisy label slices. The correlation slices are generated from CelebA (Liu et al., 2015), a facial dataset with abundant binary facial attributes like whether the person wears lipstick. Correlation slices include 520 settings. They bear resemblance to subpopulation shift (Yang et al., 2023), where a subgroup is predefined as the minor group by generating spurious correlations between two attributes when sampling training data. That subgroup also tends to be the underperforming group after training. The other two types of slices are generated from ImageNet (Deng et al., 2009), which has a hierarchical class structure. Rare slices include 118 settings constructed by controlling the proportion of a predefined subclass to be small. Noisy label slices include 248 settings formulated by adding label noise to a predefined subclass. Although many settings comprise more than one predefined error slice, we check and find that the given model actually achieves even better performance on a number of slices than the corresponding overall performance. For accurate and convenient evaluation, we select the worst-performing slice of the given model for each setting.
A.12 Examples from CivilComments
@@
When do see the piece on the worlds most annoying Catholics? Buddhist? Muslims? 375375
@@ Appendix B Related Work
It is widely acknowledged that models tend to make systematic mistakes on some subpopulations, which leads to the problem of subpopulation shift (Yang et al., 2023; Yu et al., 2024). To guarantee the worst subpopulation performance, some generate pseudo environment labels (Creager et al., 2021; Nam et al., 2021) and then apply existing invariant learning methods (Arjovsky et al., 2019; Krueger et al., 2021). Others take advantage of importance weighting to upweight the minority group or worst group like GroupDRO (Sagawa et al., 2019) and JTT (Liu et al., 2021a), or to combine with Mixup (Zhang et al., 2018) for more benefits (Han et al., 2022). A more recent work points out that current methods for subpopulation shifts heavily rely on the availability of group labels during model selection, and even simple data balancing techniques can achieve competitive performance (Idrissi et al., 2022). For better comparisons between algorithms and promotion of future algorithm development, Yang et al. (Yang et al., 2023) establish a comprehensive benchmark across various types of datasets.
@@
Instead of developing algorithms to improve subpopulation robustness (Liu et al., 2021b), the operation of error slice discovery has also attracted much attention recently. It is more flexible in that it can be followed by either non-algorithmic interventions like collecting more data for error slices, or algorithmic interventions like upweighting data belonging to error slices. There are mainly two paradigms for the process of error slice discovery. The first paradigm, also the more traditional practice, separates error slice discovery and later interpretation via case analyses or with the help of multi-modal models. Spotlight (d’Eon et al., 2022) attempts to learn a centroid in the representation space and employ the distance to this centroid as the error degree. InfEmbed (Wang et al., 2023b) employs the influence of training samples on each test sample as embeddings used for clustering. PlaneSpot (Plumb et al., 2023) concatenates model prediction probability with dimension-reduced representation for clustering. All of them interpret the identified slices via case analyses directly. Meanwhile, the error-aware Gaussian mixture algorithm Domino (Eyuboglu et al., 2022) is followed by finding the best match between candidate text descriptions and the discovered slice in the representation space of multi-modal models like CLIP (Radford et al., 2021). This paradigm has a relatively high requirement for the coherence of identified error slices so that they can be interpreted. The second paradigm incorporates the discovery and interpretation of error slices together. Both HiBug (Chen et al., 2023) and PRIME (Rezaei et al., 2024) divide the whole population of data into subgroups through proposing appropriate attributes and conducting zero-shot classification for these attributes using pretrained multi-modal models, and then directly calculate average performance for subgroups to identify the risky ones. The obtained subgroups are naturally interpretable via the combination of the attribute pseudo labels. Two recent works (Wiles et al., 2022; Gao et al., 2023) generate data from diffusion models (Rombach et al., 2022) before identifying error slices, avoiding the requirement of an extra validation dataset for slice discovery. It is obvious that the second paradigm heavily relies on the quality of proposed attributes and the capability of pretrained multi-modal models.
@@
Another branch of works sharing a similar goal with error slice discovery is error prediction (or performance prediction). Although they are also able to find slices with high error, they focus on predicting the overall error rate given an unlabeled test dataset, and measure the effectiveness of error prediction methods via the gap between the predicted performance and the ground-truth one. Moreover, they do not emphasize the coherence and interpretability of error slices. Currently, there are several ways for error prediction. Some employ model output properties on the given test data like model confidence (Garg et al., 2021; Guillory et al., 2021), neighborhood smoothness (Ng et al., 2022), prediction dispersity (Deng et al., 2023), invariance under transformations (Deng et al., 2021), etc. Inspired by domain adaptation (Long et al., 2015; Ben-David et al., 2010), some make use of distribution discrepancy between training data and unlabeled test data (Deng & Zheng, 2021; Yu et al., 2022; Lu et al., 2023). Others utilize model disagreement between two models identically trained except random initialization and batch order during training (Jiang et al., 2021; Baek et al., 2022; Chen et al., 2021; Kirsch & Gal, 2022), which exhibits SOTA performance in the error prediction task (Trivedi et al., 2023).

### v2

@@ A.4 Experimental details and related analyses
Choice of feature extractors
Following previous works of error slice discovery (Eyuboglu et al. 2022; Wang et al. 2023), for image data, we employ the image encoder of CLIP with a backbone of ViT-B/32 to extract embeddings of images for error slice discovery algorithms in our main experiments. For text data, we employ pretrained BERTbase to extract embeddings. To show that our method is flexible in the choice of the feature extractor $h_{fe}$, we conduct additional experiments on CelebA by changing CLIP-ViT-B/32 to CLIP-ResNet50 and ImageNet-supervised-pretrained ResNet50. Table 7 shows that whatever the pretrained feature extractor is, MCSD consistently identifies slices of low accuracy and outperforms other methods in terms of manifold compactness. It is worth noting that this conclusion is valid even for MCSD with ResNet50, which is generally considered as a weaker pretrained feature extractor than ViT-B/32 employed by baselines. In Figure 13, we can see that MCSD with different pretrained feature extractors truly identifies coherent error slices for the blond hair category of CelebA. As for the practice of using pretrained feature extractors, it is acceptable and generally adopted in previous works of error slice discovery (Eyuboglu et al. 2022; Jain et al. 2023; Plumb et al. 2023).
@@
[…]
@@
A.5 Time comparison
@@
A.6 Hyperparameter selection and analyses
Hyperparameter selection
@@ A.6 Hyperparameter selection and analyses
Hyperparameter analyses
In this part, we conduct hyperparameter analyses on the category “Blond Hair” of CelebA for the coherence coefficient $\lambda$, the size $\alpha$, and the number of neighbors $k$ when building the kNN graph. From Table 9, we find that both the accuracy and manifold compactness are best when $\lambda$ and $\alpha$ are in a moderate range, neither too large nor too small. This implies the importance of the balance between pursuing high error and high coherence, which could be achieved by the tuning strategy mentioned in Section A.6. This also implies the importance of appropriately controlling $\alpha$, i.e. the size of the slice, which is set according to experience in our implementation. Its selection is left for future work.
@@
[…]
@@
A.7 Performance improvement via utilization of the discovered error slices
We conduct experiments to show performance improvement that the identified error slices could bring via data collection guided by the interpretable characteristics of identified error slices, following the practice of non-algorithmic interventions of Liu et al. (2023b). For example, for a given trained image classification model on CelebA, the identified error slice exhibits characteristics of blond hair male, then we could be guided to collect specific data of the targeted characteristics of blond hair male and add to the training data, which is a non-algorithmic intervention and a straightforward and practical way of improving performance of the original model after interpreting characteristics of the identified error slice. Here we compare the results of guided data collection and random data collection. To simulate the guided data collection process, for CelebA, since the identified error slice for the category of blond hair is male and there are extra annotations of sex, we add the images annotated as blond hair male in validation data to training data. Since the identified error slice for not blond hair category is female bearing vintage styles, and there are no related attribute annotations, we directly add the images of the identified slice to the training data. For CheXpert, the identified error slices are from the frontal view for ill patients and from the left lateral view for healthy patients, and CheXpert has annotations of views, so we add the corresponding images in validation data to training data. To simulate the random data collection process, we randomly sample the same number of images from validation data and add to training data for each dataset. Then we retrain the model three times with varying random seeds.
@@ A.7 Performance improvement via utilization of the discovered error slices
Here worst group accuracy is defined following a distribution shift benchmark (Yang et al. 2023), where CelebA and CheXpert are divided into groups according to annotated attributes, and worst group accuracy is an important metric. From Table 10, we can see that guided data collection outperforms the original model and random data collection in both metrics, especially in worst group accuracy. This illustrates that our method is beneficial to performance improvement in practical applications.
A.8 Benchmark details
Dcbench (Eyuboglu et al. 2022) offers a large number of settings for the task of error slice discovery. Each setting consists of a trained ResNet-18 (He et al. 2016), a validation dataset and a test dataset, both with labels of predefined underperforming slices. The validation dataset and its error slice labels are taken as the input of slice discovery methods, while the test dataset and its error slice labels are used for evaluation.
There are 886 settings publicly available in the official repository of dcbench [^1] https://github.com/data-centric-ai/dcbench, comprising three types of slices: correlation slices, rare slices, and noisy label slices. The correlation slices are generated from CelebA (Liu et al. 2015), a facial dataset with abundant binary facial attributes like whether the person wears lipstick. Correlation slices include 520 settings. They bear resemblance to subpopulation shift (Yang et al. 2023), where a subgroup is predefined as the minor group by generating spurious correlations between two attributes when sampling training data. That subgroup also tends to be the underperforming group after training. The other two types of slices are generated from ImageNet (Deng et al. 2009), which has a hierarchical class structure. Rare slices include 118 settings constructed by controlling the proportion of a predefined subclass to be small. Noisy label slices include 248 settings formulated by adding label noise to a predefined subclass. Although many settings comprise more than one predefined error slice, we check and find that the given model actually achieves even better performance on a number of slices than the corresponding overall performance. For accurate and convenient evaluation, we select the worst-performing slice of the given model for each setting.
A.9 Examples from CivilComments
@@
When do see the piece on the worlds most annoying Catholics? Buddhist? Muslims? $375$
A.10 Code
Code is available at https://github.com/h-yu16/MCSD.
A.11 More examples for case studies
Figure 14: More examples of the category “Blond Hair” of CelebA.
Figure 15: More examples of the category “Not Blond Hair” of CelebA.
Figure 16: More examples of the category “ill” of CheXpert.
Figure 17: More examples of the category “healthy” of CheXpert.
@@ Appendix B Related Work
It is widely acknowledged that models tend to make systematic mistakes on some subpopulations, which leads to the problem of subpopulation shift (Yang et al. 2023). To guarantee the worst subpopulation performance, some generate pseudo environment labels (Creager, Jacobsen, and Zemel 2021; Nam et al. 2021) and then apply existing invariant learning methods (Arjovsky et al. 2019; Krueger et al. 2021). Others take advantage of importance weighting to upweight the minority group or worst group like GroupDRO (Sagawa et al. 2019) and JTT (Liu et al. 2021a), or to combine with Mixup (Zhang et al. 2018) for more benefits (Han et al. 2022). A more recent work points out that current methods for subpopulation shifts heavily rely on the availability of group labels during model selection, and even simple data balancing techniques can achieve competitive performance (Idrissi et al. 2022). For better comparisons between algorithms and promotion of future algorithm development, Yang et al. (Yang et al. 2023) establish a comprehensive benchmark across various types of datasets.
OOD Generalization
As a superset of subpopulation shift, Out-of-Distribution (OOD) generalization aims to address distribution shift more generally. There exists multiple branches of works. Domain generalization (Huang et al. 2020; Zhou et al. 2021; Zhang et al. 2023b, a) and invariant learning (Liu et al. 2021b, 2025, 2024; Creager, Jacobsen, and Zemel 2021) try to learn relationships that are generalizable to unseen distributions via utilizing multiple training environments. Distributionally robust optimization (DRO) (Duchi and Namkoong 2021; Volpi et al. 2018; Sinha, Namkoong, and Duchi 2018) aims to optimize the worst distribution centered around the training distribution. Stable learning (Kuang et al. 2020; Shen et al. 2020; Yu et al. 2023, 2025a) decorrelates covariates to mitigate the usage of spurious correlations.
@@
Instead of designing algorithms to improve generalization, error slice discovery has also attracted much attention recently. It is more flexible in that it can be followed by either non-algorithmic interventions like collecting more data for error slices, or algorithmic interventions like upweighting data belonging to error slices. There are mainly two paradigms for the process of error slice discovery. The first paradigm, also the more traditional practice, separates error slice discovery and later interpretation via case analyses or with the help of multi-modal models. Spotlight (d’Eon et al. 2022) attempts to learn a centroid in the representation space and employ the distance to this centroid as the error degree. InfEmbed (Wang et al. 2023) employs the influence of training samples on each test sample as embeddings used for clustering. PlaneSpot (Plumb et al. 2023) concatenates model prediction probability with dimension-reduced representation for clustering. All of them interpret the identified slices via case analyses directly. Meanwhile, the error-aware Gaussian mixture algorithm Domino (Eyuboglu et al. 2022) is followed by finding the best match between candidate text descriptions and the discovered slice in the representation space of multi-modal models like CLIP (Radford et al. 2021). This paradigm has a relatively high requirement for the coherence of identified error slices so that they can be interpreted. The second paradigm incorporates the discovery and interpretation of error slices together. HiBug (Chen, Li, and Xu 2023) and PRIME (Rezaei et al. 2024) divide the whole population of data into subgroups through proposing appropriate attributes and conducting zero-shot classification for these attributes using pretrained multi-modal models, and then directly calculate average performance for subgroups to identify the risky ones. The obtained subgroups are naturally interpretable via the combination of the attribute pseudo labels. Two recent works (Wiles, Albuquerque, and Gowal 2022; Gao et al. 2023) generate data from diffusion models (Rombach et al. 2022) before identifying error slices, avoiding the need of an extra validation dataset for slice discovery. It is obvious that the second paradigm heavily relies on the quality of proposed attributes and the capability of pretrained multi-modal models. Besides, it is worth noting that error slice discovery is closely related to the evaluation of OOD generalization (Wu et al. 2024; Yu et al. 2024a; Blanchet et al. 2024; Yu et al. 2024b, 2025b) since it is equivalent to evaluating models under subpopulation shift.
@@
Another branch of works sharing a similar goal with error slice discovery is error prediction (or performance prediction). Although they are also able to find slices with high error, they focus on predicting the overall error rate given an unlabeled test dataset, and measure the effectiveness of error prediction methods via the gap between the predicted performance and the ground-truth one. Moreover, they do not emphasize the coherence and interpretability of error slices. Currently, there are several ways for error prediction. Some employ model output properties on the given test data like model confidence (Garg et al. 2021; Guillory et al. 2021), neighborhood smoothness (Ng et al. 2022), prediction dispersity (Deng et al. 2023), invariance under transformations (Deng, Gould, and Zheng 2021), etc. Inspired by domain adaptation (Long et al. 2015; Ben-David et al. 2010), some make use of distribution discrepancy between training data and unlabeled test data (Deng and Zheng 2021; Yu et al. 2022; Lu et al. 2023). Others utilize model disagreement between two models identically trained except random initialization and batch order during training (Jiang et al. 2021; Baek et al. 2022; Chen et al. 2021; Kirsch and Gal 2022), which exhibits SOTA performance in the error prediction task (Trivedi, Koutra, and Thiagarajan 2023).